OpenStudy (anonymous):

A boat requires four hours to travel a given distance upstream against a five mile per hour current, but only requires three hours to travel downstream with that same current back to the dock. Find the speed of the boat in still water. Show work

OpenStudy (anonymous):

d=4(x-5) d=3(x+5) 4(x-5)=3(x+5) solve for x

OpenStudy (anonymous):

4x-20=3x+15 x=15+20=35